OpenStudy (anonymous):

Follow the directions for each problem to write a quadratic equation that has the given number of solutions. Be sure to show all the work leading to your answer. Think of another quadratic equation that has two (2) real number solutions. Write the equation in ax^2 + bx + c = 0 form. Then find the value of the discriminant to support your answer.

OpenStudy (anonymous):

@4everaddicted2anime

OpenStudy (4everaddicted2anime):

Here is a quadratic equation that has 2 solutions: \[x^2+3x-4=0\] Do you know how to find the discriminant?

OpenStudy (anonymous):

not off the top of my head

OpenStudy (4everaddicted2anime):

Here is the discriminant formula: \[b^2-4ac\]

OpenStudy (anonymous):

ok

OpenStudy (4everaddicted2anime):

What are the values for a b and c in the equation I gave you?

OpenStudy (anonymous):

= - 4ac + b^2

OpenStudy (4everaddicted2anime):

Quadratic equations look like this: ax^2 + bx + c =0 The equation I gave you is: x^2 + 3x -4 =0 What is the value of a? b? c?

OpenStudy (4everaddicted2anime):

a=1 b=3 c=-4 Now plug that into the discriminant formula

OpenStudy (anonymous):

ok 1^2+3x−4=0

OpenStudy (4everaddicted2anime):

no. What you should have after plugging these numbers into the iscriminant formula is: 3^2 -4(1)(-4)=?

OpenStudy (anonymous):

25

OpenStudy (4everaddicted2anime):

Good job. When the discriminant is negative there are no real solutions. When the discriminant equals zero it has one solution. When the discriminant is bigger than zero it has 2 solutions.